### v1

# Explaining word embeddings with perfect fidelity: Case study in research impact prediction

Abstract
Best performing approaches for scholarly document quality prediction are based on embedding models, which do not allow direct explanation of classifiers as distinct words no longer correspond to the input features for model training. Although model-agnostic explanation methods such as Local interpretable model-agnostic explanations (LIME) can be applied, these produce results with questionable correspondence to the ML model. We introduce a new feature importance method, Self-model Rated Entities (SMER), for logistic regression-based classification models trained on word embeddings. We show that SMER has theoretically perfect fidelity with the explained model, as its prediction corresponds exactly to the average of predictions for individual words in the text. SMER allows us to reliably determine which words or entities positively contribute to predicting impactful articles. Quantitative and qualitative evaluation is performed through five diverse experiments conducted on 50.000 research papers from the CORD-19 corpus. Through an AOPC curve analysis, we experimentally demonstrate that SMER produces better explanations than LIME for logistic regression.
keywords:
feature importance, embeddings, scientometrics
Introduction
A recent influential study by Tshitoyan et al. pointed to the ability of word embeddings to predict new thermoelectric materials that were indeed reported in future years in material engineering journals [1].
This demonstrates that embedding-based machine learning models can establish a significant relationship between certain concepts well in advance of them being recognized in a scientific field and beyond. We build on this property of word embeddings to capture latent meaning, extending the work of Tshitoyan et al., who relied on vector similarity to find novel, desirable chemical materials, to the supervised setting of building predictive machine learning models.
The emerging generation of methods for the prediction of the impact or quality of research builds upon the latest advances in text representation [2, 3, 4]. These algorithms are typically based on the projection of text elements into $\mathbb{R}^{d}$, so-called embeddings. Embeddings are real-valued vectors representing words, sentences, or whole documents that are supposed to encode semantic meaning. It is expected that word-, sentence- or document-vectors that are closer in $\mathbb{R}^{d}$ are similar in meaning.
Word embeddings perform well in tasks related to the identification of impactful research. This included the task of citation prediction [4] as well as predicting whether a paper will be accepted or rejected [3, 2]. However, the evaluations of these approaches have focused largely on model predictive performance, while many applications require reliable explanations. Having the information that a draft of the paper will not be accepted is not as useful to the author as a specific and reliable explanation of what specifically contributes to its positive or negative perception by the model. We investigated the use of several state-of-the-art model-agnostic methods for explanations, including LIME and Shapley plots, in our previous work [5], noting that the quality of the explanations generated for why specific research was cited lags behind white-box models such as rule learning. The disadvantage of these methods is that they obscure the explanation process by introducing randomly generated data and additional machine-learning models. The quality of such explanations is not yet thoroughly studied, with research showing that methods such as LIME can be manipulated to have very low fidelity [6].
In this article, we, therefore, propose a new explanation method that provides intrinsic explanations of embedding-based predictive models without training additional models or generating new instances. Furthermore, we theoretically show that the newly proposed explanation method is guaranteed to have perfect fidelity for logistic regression predictive models. In the five experiments, we confirm this result empirically.
Method
This section describes the workflow of our methodology for predicting the impactfulness of words as well as entire documents. A reliable prediction of research impact for any research paper only based on its abstract is a difficult task. In our research, we address a narrower subproblem of predicting which research will be groundbreaking, ie., we are interested in identifying the top most influential research – both articles and newly emerged entities, such as drugs or virus strains in the CORD-19 training data context. The focus on the most confident predictions makes the research impact and explainability problem more tractable.
We first describe our new method of building an embedding-based classifier for predicting article impact, which will directly provide word-based explanations. Then, we will cover two existing baseline methods (ORC and LIME). Based on this, we will provide an overview of the framework for predicting the impact of research articles and impactful entities (words). Finally, we will cover the methods we used for evaluation in the five experiments.
Self Model Entities Rated (SMER)
We represent a scientific article by its abstract, which is a string consisting of dictionary words. Let $\mathcal{W}$ stand for a dictionary; an element $w\in\mathcal{W}$ is called a word. A document, a text—or specifically in this study, an abstract of a scientific paper—is a nonempty sequence $A=(w_{1},\dots,w_{N})$ of words. We are given a family $\mathcal{A}$ of abstracts $A_{i}=(w_{i,1},\dots,w_{i,N_{i}})$, $i=1,\dots,n$, called corpus. We assume that $\mathcal{W}=\bigcup_{i=1}^{n}\bigcup_{j=1}^{N_{i}}w_{i,j}$.
Given a text corpus as a training set, the trained model assigns a score to each text feature. Thus the score estimated by our model provides a quantitative measure of feature importance.
Embedding of words
We use the corpus $\mathcal{A}$ to produce an embedding $\widetilde{w}\in\mathbb{R}^{d}$ for each word $w\in\mathcal{W}$. In this work, we assume the use of non-contextual embeddings, which represent the dictionary $\mathcal{W}$ as a family $\{\widetilde{w}\in\mathbb{R}^{d}\mid\ w\in\mathcal{W}\}$ of points in $\mathbb{R}^{d}$ which can be subsequently used as inputs to other procedures requiring real-valued input data.
The most significant property of embeddings for this study is that when words $w,w^{\prime}\in\mathcal{W}$ frequently appear “close” to each other in the training dataset $\mathcal{A}$, then $D(\widetilde{w},\widetilde{w}^{\prime})$ is “small”, where $D$ stands for some distance measure on $\mathbb{R}^{d}$, such as the cosine similarity or Euclidian distance.
Embedding of documents
Recall that an abstract $A$ of a scientific paper is here formalized simply as a sequence of some dictionary words $w_{1},\dots,w_{N}$, possibly with repetitions. The aim is to construct $\widetilde{A}\in\mathbb{R}^{d}$, an embedding of $A$, a representation of abstract as a vector in $\mathbb{R}^{d}$. The function $A\mapsto\widetilde{A}$ should fulfil certain natural properties, such as
  1. 1.
    (reflexivity) if $A=(w,w,\dots,w)$, then $\widetilde{A}=\widetilde{w}$;
  2. 2.
    (replacement principle) if $A=(w_{1},\dots,w_{j-1},w_{j},w_{j+1},\dots,w_{N})$,
    $A^{\prime}=(w_{1},\dots,w_{j-1},w_{j}^{\prime},$ $w_{j+1},\dots,w_{N})$ and $\|\widetilde{w}_{j}-\widetilde{w}_{j}^{\prime}\|\leq\varepsilon$, then $\|\widetilde{A}-\widetilde{A}^{\prime}\|\leq\varepsilon/N$;
  3. 3.
    (extension principle) if $A=(w_{1},\dots,w_{j},w_{j+1},\dots,w_{N})$, $A^{\prime}=(w_{1},\dots,w_{j},\\ w^{\prime},w_{j+1},\dots,w_{N})$ and $\|w^{\prime}-w_{\iota}\|\leq\varepsilon$ for every ${\iota}=1,\dots,N$, then $\|\widetilde{A}-\widetilde{A}^{\prime}\|\leq\varepsilon/N$.
Reflexivity states that if an abstract $A$ contains a single dictionary word (possibly with multiple occurrences), then the embedding of that abstract coincides with the embedding of the word. The replacement principle is a form of continuity: if we replace a word $w_{j}$ in an abstract with a close word $w_{j}^{\prime}$, then the embedding of the abstract cannot change too much. Finally, the extension principle is another form of continuity: if we have an abstract $A$ and we extend it by adding a new word $w^{\prime}$, which is close to all words $w_{1},\dots,w_{N}$, then the embedding of the extended abstract $A^{\prime}$ cannot be too far from $A$.
In this study, we aggregate embeddings of words appearing in an abstract as follows: if $A=(w_{1},\dots,w_{N})$, then
$\widetilde{A}=\frac{1}{N}\sum_{j=1}^{N}\widetilde{w}_{j},$ (1)
which we refer to as weighted-average embedding. (The average is weighted in the sense that a word is counted in the above sum as many times as it appears in $A$.) For dataset $\mathcal{A}$, we have $\widetilde{A}_{i}=N_{i}^{-1}\sum_{j=1}^{N_{i}}\widetilde{w}_{i,j}$, $i=1,\dots,n$.
Remark
It is easy to see that (Eq. 1) fulfills properties 1–3. Reflexivity is obvious. The replacement principle follows from the fact that $\|\widetilde{A}-\widetilde{A}^{\prime}\|=\|\frac{1}{N}\sum_{\iota=1}^{N}\widetilde{w}_{\iota}-\frac{1}{N}[\widetilde{w}_{j}^{\prime}+\sum_{\iota\in\{1,\dots,N\}\setminus\{j\}}\widetilde{w}_{\iota}]\|=\|\frac{1}{N}(\widetilde{w}_{j}-\widetilde{w}_{j}^{\prime})\|\leq\varepsilon/N$. The extension principle is proved analogously: $\|\widetilde{A}-\widetilde{A}^{\prime}\|=\|\frac{1}{N}\sum_{\iota=1}^{N}\widetilde{w}_{\iota}-\frac{1}{N+1}[\widetilde{w}^{\prime}+\sum_{\iota=1}^{N}\widetilde{w}_{\iota}]\|=\|\frac{1}{N(N+1)}\sum_{\iota=1}^{N}\widetilde{w}_{\iota}-\frac{N}{N(N+1)}\widetilde{w}^{\prime}\|=\|\frac{1}{N(N+1)}\sum_{\iota=1}^{N}(\widetilde{w}_{\iota}-\widetilde{w}^{\prime})\|\leq\|\frac{1}{N(N+1)}N\varepsilon\|=\varepsilon/(N+1)\leq\varepsilon/N.$
A classifier: Logistic regression
The next step is the training of a classification model. Assume that a target variable is available for supervised learning. That is, let our dataset consist of $(A_{i},y_{i})_{i=1,\dots,n}$, where $y_{i}\in\{0,1\}$ is a measure of a research impact of a paper with abstract $A_{i}$. In our work, values $\{0,1\}$ are used as indicators of whether the research is highly cited/impactful (coded as 1), or not (coded as 0).
Logistic regression is used here as a natural representative of classifiers. Given that it is a linear additive model, it is often used as a base learner when an explanation is needed [7]. Recall that logistic regression assumes the regression relationship
$\mathsf{E}[y_{i}\mid A_{i}]=L(\beta_{0}+\beta^{T}\widetilde{A}_{i}),\quad i=1,\dots,n,$
with
$L(\xi)=\frac{1}{1+e^{-\xi}}$ (2)
being the logistic function and $\beta_{0}\in\mathbb{R}$, $\beta\in\mathbb{R}^{d}$ parameters. (Here, $\mathsf{E}[\,\cdot\!\mid\!\cdot\,]$ stands for the conditional expectation.) Training the model amounts to finding estimates $\widehat{\beta}_{0},\widehat{\beta}$ of the model parameters $\beta_{0},\beta$ e.g. by minimization of the residual squares $\sum_{i=1}^{n}[y_{i}-L(\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{A}_{i})]^{2}$ or by maximization of the log-likelihood $\sum_{i=1}^{n}\left[y_{i}\log L(\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{A}_{i})+(1-y_{i})\log(1-L(\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{A}_{i}))\right]$, often with restrictions on the norm $\|(\beta_{0},\beta^{T})\|$ (e.g. by adding a constraint into the optimization or by adding a regularization penalty into the objective function).
Score of an abstract
Given an abstract $A$, we can predict its impactfulnessusing the trained logistic model by computing
$\textit{SCORE}(A):=L(\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{A}).$ (3)
SMER Score: Logistic regression on word embeddings is self-explainable
The basis of our self-explanatory embeddings-based classification framework is the observation that a word $w\in\mathcal{W}$ can be seen as a single-word abstract $A=(w)$. We can thus compute feature importance (i.e. word importance) directly using the same model, which was trained to classify all documents (the SCORE from Eq. 3). When applied to a single word, the formula reduces to
$\textit{SCORE}(w)=L(\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{w}).$ (4)
Words $w\in\mathcal{W}$ can be sorted according to their scores; by selection of a cutoff threshold, we get a classifier of high- and low-scoring words, i.e., words that have high or low potential to contribute to the number of article’s citations. Since $\widetilde{w}$ is an embedding vector, the method can also score words unseen in the training data as long as these can be assigned an embedding vector.
Properties of SMER
Notation
For the sake of brevity, let us introduce abbreviations for the logits
$\lambda(w):=\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{w},\quad\lambda(A):=\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{A},$
where $w$ is a word and $A$ is an abstract.
Recalling that $L$ is the logistic function (2), we have $\textit{SCORE}(w)=L(\lambda(w))$, $\textit{SCORE}(A)=L(\lambda(A))$. The following observation shows an important property.
Proposition 1 (linearity of logits)
If $A=(w_{1},\dots,w_{N})$ is an abstract, then
$\lambda(A)=\frac{1}{N}\sum_{\ell=1}^{N}\lambda(w_{\ell}).$
Proof. Using (1) we can write
$\lambda(A)=\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{A}=\widehat{\beta}_{0}+\widehat{\beta}^{T}\left(\frac{1}{N}\sum_{\ell=1}^{N}\widetilde{w}_{\ell}\right)=\frac{1}{N}\sum_{\ell=1}^{N}\left(\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{w}_{\ell}\right)=\frac{1}{N}\sum_{\ell=1}^{N}\lambda(w_{\ell}).\qed$
Proposition 1 is a key property for assessing the fidelity of SMER.
Fidelity is defined as the percentage of predictions that are the same between the original model and the explanation [8, 9]. Proposition 1 tells us that the average of the individual explanations (predictions for word vectors of individual words) is equal to the prediction of the explained model (the document classifier). As a consequence, the SMER explanation technique has 100% fidelity.
Proposition 2 (preservation of ordering)
The ordering of dictionary words according to their SCOREs is the same as their ordering according to the logits ($\lambda(w)$-values).
Proof. Since $\textit{SCORE}(w)=L(\lambda(w))$ and $L$ is an increasing function, for any pair of dictionary words $w,w^{\prime}$ we have $\textit{SCORE}(w)\leq\textit{SCORE}(w^{\prime})$ if and only if $\lambda(w)\leq\lambda(w^{\prime})$. ∎
Proposition 2 tells us that if we want to sort the dictionary words according to their scores, we can equivalently sort them according to their logits ($\lambda(w)$-values). However, in contrast to the scores, the logits do not possess the nice property of lying in the interval (0,1).
Proposition 3 (monotonicity of SCOREs)
If $A=(w_{1},\dots,w_{k-1},w_{k},w_{k+1},\dots,w_{N})$ and $A^{\prime}=(w_{1},\dots,w_{k-1},w^{\prime}_{k},w_{k+1},\dots,w_{N})$ are abstracts, then
$\textit{SCORE}(w^{\prime}_{k})>\textit{SCORE}(w_{k})\ \ \Longrightarrow\ \ \textit{SCORE}(A^{\prime})>\textit{SCORE}(A).$
(In other words: if the score of a word in an abstract increases, then the score of the abstract increases as well.)
Proof. $L$ is an increasing function; thus $\textit{SCORE}(w^{\prime}_{k})>\textit{SCORE}(w_{k})$ implies $\Delta:=\lambda(w^{\prime}_{k})-\lambda(w_{k})>0$. Now
$\textit{SCORE}(A)$ $=L(w(A))=L\left(\frac{1}{N}\sum_{\ell=1}^{N}\lambda(w_{\ell})\right)<L\left(\frac{1}{N}\sum_{\ell=1}^{N}\lambda(w_{\ell})+\frac{\Delta}{N}\right)$
$=L\left(\frac{1}{N}\left(\sum_{\ell\in\{1,\dots,N\}\setminus\{k\}}\lambda(w_{\ell})\right)+\frac{\lambda(w^{\prime}_{k})}{N}\right)=L(w(A^{\prime}))=\textit{SCORE}(A^{\prime}).$
The strict inequality again follows from the fact that $L$ is increasing.∎
Reference baseline feature importance: ORC and LIME
In this section, we define two reference models. One is based on training a logistic regression model over bag-of-words, which is the previous-generation approach for representing document content. This method is also intrinsically explainable as the importance of individual words follows from their assigned logistic regression coefficients. Finally, we cover LIME, a state-of-the-art technique commonly used to explain embedding-based classification models.
BoW representation of abstracts
We continue with the notation from the previous section. Let the dictionary $\mathcal{W}$ consist of $M$ words $v_{1},\dots,v_{M}$. Bag-of-Words (BoW) represents an abstract $A$ by a vector $\bar{A}\in\mathbb{R}^{M}$ of dictionary words present in the abstract:
$\bar{A}_{k}=\left\{\begin{array}[]{ll}1,&\text{\ \ \ if $v_{k}\in A$,}\\ 0&\text{\ \ \ otherwise,}\end{array}\right.\quad\quad k=1,\dots,M.$ (5)
Logistic regression
The corpus $\mathcal{A}$ is represented by vectors $\bar{A}_{1},\dots,\bar{A}_{n}\in\mathbb{R}^{M}$ which serve as explanatory variables in the logistic regression model
$\mathsf{E}[y_{i}\mid A_{i}]=L(\gamma_{0}+\gamma^{T}\bar{A}_{i}),\quad i=1,\dots,n,$
with coefficients $\gamma_{0}\in\mathbb{R}$ and $\gamma\in\mathbb{R}^{M}$. The estimate of $\gamma$ is denoted by $\widehat{\gamma}$.
ORC: Odds ratio change as a measure of word importance in BoW models
Observe that $\widehat{\gamma}_{k}>0$ indicates that adding the word $v_{k}$ into an abstract (if not yet present) increases the odds of the article becoming highly cited. On the other hand, $\widehat{\gamma}_{k}<0$ indicates that adding $v_{k}$ into an abstract is predicted to decrease the odds of the article becoming highly cited. Therefore, the coefficient $\widehat{\gamma}_{k}$ measures the impact of the dictionary word $v_{k}$ on the target variable $y$. Recall that within the logistic regression framework this effect can be interpreted in terms of odds ratio. We define
$\textit{ORC}(v_{k}):=e^{\widehat{\gamma}_{k}},\quad k=1,\dots,M.$ (6)
For BoW, entity (word) importance thus directly corresponds to feature importance as measured by Odds Ratio Change (ORC), because features are binary word indicators.
LIME: Locally Interpretable Model Explanations
The LIME method (Local Interpretable Model-agnostic Explanations) [10] is a technique for explaining model decisions at the level of individual data instances. This method focuses on creating an interpretable model that approximates the behavior of the original model in the vicinity of a specific data instance.
The explanation formulated by the LIME method for a word $v_{k}$ is expressed as:
$E(v_{k})=L(f,g,\Pi_{v_{k}})+\Omega(g),$
where $E(v_{k})$ represents the explanation for the word $v_{k}$, $L(f,g,\Pi_{v_{k}})$ denotes the measurement of unfaithfulness of the model $g$ when approximating $f$ in the locality $\Pi_{v_{k}}$, and $\Omega(g)$ represents the complexity of the model $g$.
Framework for predicting impactful research
Our empirical study employs the methodology from Figure 1. Here, we cover the technical details of the phases depicted in the figure: feature engineering, target derivation, training of the classification model, estimating feature importance, model validation, and domain specification.
Figure 1: Methodological pipeline of workflow steps.
Data sources
The data utilized for this study is derived from the CORD-19 corpus [11], which consists primarily of research articles pertaining to biology and medicine relevant to SARS-CoV-2 and that were published between 1952-2023. Specifically, we employed the version of the corpus downloaded on April 23, 2022, comprising a total of 992,921 articles. Article citation counts were extracted from Opencitations.org.
For the training of classifiers, we used only article abstracts from the years 2016-2021 for which the OpenCitations data were available. The data sizes used for individual experiments are summarized in Table 1. As the table shows, more article abstracts were available for the training of embeddings because, for this task, labels derived from OpenCitations were not required.
Publication years Used in Count of abstracts for training of
Word embeddings Classification
2016+2017 Experiment 1 12,589 1,867
2017+2018 Experiment 1 13,036 1,916
2018+2019 Experiment 1 14,503 2,256
2019+2020 Experiment 1 + 2 +3 132,422 18,571
2020+2021 Experiment 4 + 5 414,174 6,964
Table 1: Number of abstracts used in training embedding models and classification models.
Preprocessing
The number of articles was reduced to 598,587 after filtering out abstracts with duplicate document object identifiers (DOI). Subsequently, articles whose number of characters did not exceed 10 were removed, reducing the article count to 483,781 articles. Filtering of non-English abstracts detected using the pre-trained Fasttext [12, 13, 14] language identification model reduced the count to 477,237 articles. For training, only articles with recent years of publication (2016-2021) were used, which can be seen in Table 1. For training classification models, 80% of all articles with citation data for the specified years of publications were used, which is the number shown in the last column of Table 1. The total number of unique abstracts we used for training embeddings was 225,679. Note that to train word embedding models, more articles were available as abstracts with missing citation data could also be used.
The abstracts were preprocessed as follows. Punctuation was removed except for dashes (as these are used in compound biomedical names), all text was converted to lowercase, and single numbers were deleted. Lemmatization was used to convert the words into root words. After this preprocessing, articles with less than 20 characters were removed. The final dataset contained 476,175 articles. After preprocessing, we also checked for possible duplications, determining that more than 99.8% of abstracts were unique.
To train embeddings, we used the Skip-Gram version of Word2vec, which is based on a prediction of the context of a given word. Although the alternative CBoW models are known for their speed, the Skip-Gram model is recommended for datasets with infrequent words [15], which corresponds to our focus on emerging research concepts. The skip-gram model was already used for this purpose in related work [1].
The counts of articles used to train the embeddings and the classification models differed. For the unsupervised task of embedding training, articles without available citation counts could be used, while for learning the classifier, the availability of citation counts was required to construct the target. The sizes of the respective datasets are summarized in Table 1.
Predictors
After the preprocessing steps, article abstracts were represented using two variant methods: first, embeddings-based word vectors generated from abstracts with Eq.1, and second, we used an intrinsically explainable model based on Bag of Words (see Eq. 5).
However, unlike several recent approaches, we deliberately excluded the author information and other metadata, such as journal impact factor, from training since we have found in our previous work that it can lead to publication biases [5].
The details of data used to train the embedding models for individual experiments are presented in Table 1. The dimension $d\geq 1$ is a parameter used to train the embedding models. For training the word embeddings, we used embedding vector dimensions of 100 and a window size of 10. These settings were based on our prior experience combined with recommendations from the literature [16].
Citation count as a proxy for research significance
Citation counts are a commonly used proxy for research impact obtainable on a large scale but with some limitations. For example, recent authoritative research on this topic [17] states: ”it should be strongly emphasized that citation counts correlate with the scientific relevance or impact of a scientific publication, but they do not always measure the relevance of a specific scientific publication”.
This shows that while citations often fail to reflect the true impact for various reasons, this often happens on an individual level, which is not relevant to our work as our machine learning model learns an abstraction. Our research is based on this widely held assumption, well summarized by [18]: ”Although most citation analysts seem to agree that citing or referencing is biased, it has been argued that this bias is not fatal for the use of citation as performance indicators—to a certain extent, the biases are averaged out at aggregated levels … the presence of different cognitive meanings of citations and motivations for citing does not necessarily invalidate the use of citations as (imperfect) performance measures…”.
An example of a well-documented bias is that in biosciences, positive (statistically significant) results are cited more frequently [19]. This is a potentially favourable bias for our research, as we expect impactful new concepts or research to be mostly supported by statistically significant evidence.
Citation target construction
To define the target, we combine two subtasks of scholarly document quality prediction: predicting citation counts and predicting whether a paper will be rejected or accepted [3]. We define the target variable $y_{i}$ as a 0-1 indicator of being a highly cited article. To define the notion precisely, let
$\textit{CIT}(i,t)$ $:=\text{number of citations of article $A_{i}$ originated in year$t$},$ (7)
$\textit{PUB}(i)$ $:=\text{year of publication of article $A_{i}$.}$ (8)
For example, $\textit{CIT}(i,\textit{PUB}(i))$ is the number of citations of $A_{i}$ from its publication year. By definition, if $t<\textit{PUB}(i)$, then $\textit{CIT}(i,t)=0$. The number $\sum_{t=\textit{PUB}(i)}^{2022}\textit{CIT}(i,t)$ represents the citation count of $A_{i}$ from its publication year up to the moment of our download of the citation data.
Balancing class distributions
Recall that $n$ stands for the number of corpus articles. Both training and evaluation are performed on distinct time slices defined by the set $T$ of publications years, e.g. $T=\{2016,2017\}$ and a fixed citation year $\tau$, e.g. $\tau=2018$. Then, article $A_{i}$ with $\textit{PUB}(i)\in T$ is defined to be highly cited with respect to $T$ and $\tau$ if the number of its citations from year $\tau$ (strictly) exceeds the median:
$y_{i}$ =$=\left\{\begin{array}[]{ll}1,&\text{if $\textit{CIT}(i,\tau)>\text{median}_{1\leq k\leq n}\{\textit{CIT}(k,\tau)\mid\textit{PUB}(k)\in T\}$,}\\ 0&\text{otherwise,}\\ \end{array}\right.$ (11)
where $\hskip 142.26378pt\mbox{where }i\in\{\iota\mid 1\leq\iota\leq n,\textit{PUB}(\iota)\in T\}.$ (12)
Table 2 presents the composition of the low- and highly-cited articles in the dataset after the target has been derived. The count of articles that were published in a given year differs in individual citation years due to missing citation data. These were not consistently available for each article in all covered years. Some articles are cited before they have been published — this situation can occur at the preprint stage of the article. Following the strict inequality in definition (12), the two subsets separated by the median are not completely equally sized. For example, assume a corpus containing five articles with citation counts $0,1,1,1,2$, then as a result of the preprocessing, there will be only one article in the “high” category, i.e. article with the citation count strictly above the median $1$, and the remaining four articles with at most one citation will be assigned to the “low” category.
Citation Year
2017 2018 2019 2020 2021 2022
Publication Year low high low high low high low high low high low high
2016 962 629 903 799 867 861 918 871 305 157 2 0
2017 375 368 979 616 894 874 1,057 797 341 139 1 1
2018 13 0 403 397 1,020 771 1,040 997 350 167 1 0
2019 2 1 13 1 532 497 1,284 1,248 441 192 0 0
2020 9 0 5 0 120 6 11,955 8,727 5,204 3,458 11 0
2021 0 0 0 0 0 0 53 29 33 10 0 0
Total 2,990 2,386 3,946 3,158 5,009 4,396 17,970 14,106 7,216 4,354 16 1
Table 2: Number of low- and highly-cited articles by publication and citation year in the used subset of the CORD-19 corpus used for training and evaluation of classification models.
Evaluation methodology
Data-splitting strategies
The time-based nature of the data requires a specific approach to data splitting. We report results using the time-split and random-split techniques, two commonly used approaches. Since the last year needs to be used for testing, models trained under time-split testing are not trained on the latest articles, which negatively affects their ability to evaluate the impact of recent research. We, therefore, also report results using a third technique, which addresses this issue, which we call semi-split testing.
Training set Testing set Split Type AUC – Extreme 10% of Testing set
Publication year Cit. year Publication year Cit. year BoW W2v
80% subset 2016, 2017 2017 20% subset 2016, 2017 2017 random-split 0.60 0.71 {1}
all 2016, 2017 2017 random-split 0.98 0.81 {2}
all 2018 2018 time-split 0.76 0.80 {3}
80% subset 2017, 2018 2018 20% subset 2017, 2018 2018 random-split 0.79 0.78 {4}
all 2017, 2018 2018 random-split 0.94 0.80 {5}
all 2019 2019 time-split 0.59 0.67 {6}
80% subset 2018, 2019 2019 20% subset 2018, 2019 2019 random-split 0.75 0.76 {7}
all 2018, 2019 2019 random-split 0.96 0.75 {8}
all 2020 2020 time-split 0.52 0.47 {9}
80% subset 2019, 2020 2020 20% subset 2019, 2020 2020 random-split 0.65 0.83 {10}
all 2019, 2020 2020 random-split 0.93 0.83 {11}
20% subset 2020 2021 semi-split 0.73 0.77 {12}
Table 3: Quantitative validation of LR-W2v and LR-BoW models. The best result for each evaluation setup is in bold. The time-split/semi-split is considered as most representative (in italics).
Random-split testing. Random, time-sensitive splits of data are a common evaluation technique. In our setup, 80% of data of selected periods is used to train the models and the remaining 20% data from the same periods is used for testing.
Time-split testing. This split is stricter than random-split testing, as the testing set consists of articles that were both published and cited later than the training set in order to prevent any leakage from the test data to the training data.
For training, we use 80% of available articles from publication years $t-1,t-2,\dots$ with the target being derived from citation counts of these articles accumulated by year $t-1$. The test articles were those published in year $t$ with the target being derived from citation counts also from year $t$. This is a particularly challenging setup because the prediction of citations for articles published in the same year is often very difficult.
Semi-split testing. Since the last year needs to be used for testing, models trained under time-split testing are not trained on the latest articles, which negatively affects their ability to evaluate the impact of recent research. This semi-split evaluation technique best corresponds to how we envisage that the models would be used by leveraging the latest available citation counts for testing.
Focus on groundbreaking research
To evaluate the ability of the model to separate groundbreaking research from uninfluential articles, we construct a special testing set Extreme10%. Using citation count as a proxy for the impactfulness of the research. This is a subset of the respective test set containing a selection of 5% of articles with the highest citation counts and 5% of articles with the lowest citation counts.
Evaluation measures for the prediction task
The Area Under the ROC Curve (AUC) is used as the main metric to evaluate predictive performance. Since the target is binary and both classes are balanced, the values of AUC are almost identical to the values of accuracy and F1-Score, which are thus not reported.
Evaluation scenarios
The last column of Table 3 identifies the given combination of publication and citation year used in the train and test data and type of split as the evaluation scenario. For example, for training the classification model in scenario 1 (publication years 2016 and 2017), 80% of all articles published in 2016+2017 with citations in 2017 were used, which makes a total of (962+375+629+368) = 2334 x 0.8 = 1867 articles, which also corresponds to the value in Table 1.
Evaluation of explanations through correlation of feature importance with actual impact
A key element of our research is being able to identify which words or entities (drugs, vaccines, etc.) are linked to high research impact. For a well-performing measure of feature importance, we should be able to observe a positive correlation between feature importance and the presence of the word among articles with observed research impact. High positive feature importance (for the highly cited class) will correlate with the presence of the word among highly cited papers, and vice versa for negative feature importance.
With this intuition, we define the actual impact of the word $w$ as
$\textit{ActImpact}(w)=\frac{\#\{i\in\{1,\dots,n\}\mid w\in A_{i},\ i\in H_{2021}\}}{\#\{i\in\{1,\dots,n\}\mid w\in A_{i},\ \textit{PUB}(i)\in\{2019,2020\}\}},$ (13)
where ##\## denotes the number of elements of a set and $H_{2021}$ is the set of highly-cited articles in 2021 computed as follows:
$H_{2021}$ =$=\{i\in\{1,\dots,n\}\mid PUB(i)\in\{2019,2020\},$
$\hskip 14.22636pt\,\ \textit{CIT}(i,2021)>\textit{median}_{1\leq k\leq n}\{\textit{CIT}(k,2021)\mid\textit{PUB}(k)\in\{2019,2020\}\}\},$
where functions CIT and PUB were defined in Eq. (7,8).
Given a trained model, we will use the top 100 words with the highest SCOREs (for LR-W2v) or ORCs (for LR-BoW) and bottom 100 words with the lowest values and compare these with the actual impact according to (13). The baseline for this comparison is the probability of an article being highly cited, computed as:
$p_{high}=\frac{\#H_{2021}}{\#\{i\ \mid\ \textit{PUB}(i)\in\{2019,2020\}\}}.$ (14)
Evaluation of explanations through feature perturbations
To compare the two feature importance methods used in our work, we used a standard approach based on Area Over the Perturbation Curve (AOPC) evaluation [20]. Specifically, we employed AOPCglobal [21], which is a modification of AOPC for benchmarking global feature importance measures. AOPCglobal, unlike the original AOPC, which assesses importance based on local explanations, operates by sequentially removing words based on their global importance as determined by the global feature importance. The SMER method readily provides global importance scores. To obtain these for LIME, which is a local method, we utilized GALE (Global Aggregations of Local Explanations) [21]. We used two variations of GALE: Global LIME Importance and Global Average LIME Importance [21]. Next, features were removed in the order of global importance. The removals were cumulative, first feature, first two features, etc. After each removal, the model was applied to the test set (with removed words), and the average change (typically decrease) in the predicted class probabilities was recorded. Based on this, the AOPCglobal curve was constructed by plotting the number of removed words (X-axis) against the decrease in the class probabilities (Y-axis).
To quantify the aggregate loss in predictive accuracy consequent to the systematic removal of the most significant features, we involved a trapezoidal rule integration, which was used to compute the area under the AOPCglobal curves. The curves were divided into equidistant segments, and areas of individual trapezoids were summed. For the evaluation of global metrics, the curve was segmented into 10 parts, corresponding to each curve consisting of a total of 11 data points (equating to the removal of 0 to 10 words).
Identification of the newly emerged impactful entities
The goal of the application of SMER in biosciences is to identify top candidates for important concepts, such as drugs and vaccines. The importance of a word $w$ is reflected in the value of $SCORE(w)$. To have one compound measure reflecting the degree of impact (SCORE) as well as the size of impact (NumArt), we introduce the customized score:
$\textit{CScore}(w)=\textit{SCORE}(w)\times\textit{NumArt}(w),$ (15)
where NumArt(w) is the number of articles where the word appears:
$\textit{NumArt}(w)=\#\{i\in\{1,\dots,n\}\mid w\in A_{i},\ 2020\leq\textit{PUB}(i)\leq 2022\}\}.$ (16)
Evaluation of predictions of impactful articles through bibliometric indicators
Bibliometric indicators are widely accepted to correlate with research quality. For example, combined impact factor and citation counts were previously used to predict research impact [22]. As a supplementary approach for validating predictions, we use bibliometric information related to the journal where the article appeared, as well as a bibliometric profile of its authors. We emphasize that this information has not been used in our training setup and is used only for validation.
The quality of the journal is measured through the ISI impact factor, which is computed as the citation counts to articles the journal published in 2019-2020 from articles published in the year 2021 in other journals with an impact factor divided by the number of articles published by the journal in the years 2019-2020. As a reference for a high-impact prediction, we use the median impact factor among the top 10% journals in the category “Health sciences” category (OECD category FORD 3030, data for the year 2021). To represent the research reputation of the authors, we use the maximum of the first and last authors’ Hirsch-index (H-index) values. The H-index is defined as the number of the number of articles published by the author that have at least as many citations as this number. As a reference for a high-impact prediction, we use the median H-index of full professors, which was determined by analyzing the medical education journal editorial boards [23]. Previous research has found that the H-index was a good indicator of citation in the first few years after publication [22].
We also include additional impact metrics from social media as retrieved from Altmetrics.com in September 2022: the number of tweets mentioning the research and the aggregate attention received on social media (such as the number of Facebook likes) weighted by the relative reach of each type of source (Facebook, YouTube, Twitter).
Results
Experiment 1: Predictions of impactful articles
The goal of this experiment is to evaluate the predictive performance on the task of identifying the most (and least) promising research. We compare the proposed method, logistic regression over word embeddings (further referred to as Word2Vec or w2v), with an older approach based on bag of words (BoW).
To train the models, we created four datasets, each covering articles published in two consecutive years from the period 2016-2020. The overview of the datasets is in Table 3.
For models using embeddings, for each of these datasets, embeddings were trained separately on abstracts published in the corresponding years.
The results are presented in Table 3. Although in terms of random-split evaluation, the BOW method most often has a higher AUC , LR-W2v model outperforms LR-BoW in most of the time-split & semi-split scenarios ($\{3\}$,$\{6\}$,$\{12\}$). The probable reason why in scenario $\bf\{9\}$ both models are failing (as AUC is close to 0.5) is that the training occurred on pre-pandemic data, and articles published during the pandemic were used for testing. The interest of researchers (and consequently the citations) shifted considerably as a result of the pandemic.
On the reference evaluation setup $\{12\}$, which was evaluated based on the most recent set of citations, the best AUC of 0.77 was obtained by the embeddings-based model (trained with word2vec). Overall, the embedding model seems to be better able to adapt to concept drift as measured in the time-split scenarios, where pre-pandemic data were used for training and evaluated on articles that originated subsequently.
Experiment 2: Correlation of feature importance with actual impact
This is the first of two experiments aimed at evaluating the quality of feature importance ratings provided by SMER and comparing them against a baseline. As the baseline method in this experiment, we use coefficients from logistic regression over bag-of-words. The second experiment (Experiment 3) will present the comparison against LIME.
In the BoW mode, Coefficients for individual features ($\textit{ORC}(w)$ of a word $w\in\mathcal{W}$ from Eq.6) directly correspond to word importance. As a second method, we represent the documents with uninterpretable features (word embeddings) and also train a logistic regression classifier. We then use the SMER method ($\textit{SCORE}(w)$) to identify specific words responsible for the classification. As was shown in Proposition 1, this explanation has also 100% fidelity as the average of predictions for individual word vectors is equal to the prediction of the explained model.
The logistic regression was trained on articles published in 2019-2020 for which citation data were available. The overview of the dataset is in Table 3 (scenario #12). The citation counts from 2021 were used for the validation.The models were trained with and without regularization. We chose the regularized versions (ridge regression) since they performed better in terms of correlation with ActImpact for both models.
Figure 2 answers the question of whether the top 100 words with the highest predicted values, denoted as $T_{100}^{W2V}$ and $T_{100}^{BOW}$, frequently appear in articles that later turned out to be highly cited and whether words with the lowest predicted values, denoted as $L_{100}^{BOW}$ and $L_{100}^{W2V}$, appear in abstracts of articles that later turned out to be lowly cited.
Figure 2: Evaluation of feature importance values for SMER and a baseline method using the Pearson correlation coefficient between the Actual impact of the word and its predicted feature importance. Left: The horizontal axis shows the score of a regularized Logistic Regression model trained on Word2vec embeddings (LR-W2V reg). Right: Reference method based on regularized Logistic Regression Odds Ratio Coefficients for bag of words (LR-BoW reg). Vertical axis indicates the proportion of the articles containing a given word which belongs to the highly cited category. Dots correspond to words $w\in T_{100}^{W2V}\cup L_{100}^{W2V}$ (left) and $w\in T_{100}^{BOW}\cup L_{100}^{BOW}$ (right figure). The red line corresponds to a random baseline ($p_{high}\approx 0.57$).
The SMER method has a higher correlation coefficient with the ground-truth word importance (ActualImpact) than the baseline ORC method. This shows that even though the underlying classification model uses uninterpretable features, the SMER method was able to identify specific words responsible for the classification. Note that SMER results are even better than for the ORC method, which uses a representation where individual features directly correspond to word importance.
Experiment 3: Comparing SMER vs LIME through feature perturbations
The previous two experiments showed that word embeddings are superior to the bag-of-words representation both in terms of predictive performance and, when coupled with the proposed SMER method, also when it comes to explanations. The method that we used in Experiment 2 to measure the quality of explanations was based on correlation with actual impact. In this next experiment, we compare SMER with LIME, a frequently used baseline explanation approach.
The data utilized for training the embeddings and the logistic regression model were the same as detailed in Experiment 2, which comprised articles published in years 2019 and 2020, with citation counts from 2021. For the application of both LIME and SMER explanations, we used the logistic regression model trained in Experiment 2.
LIME explainer was configured to use 15 features (with 10 being the default). As a result, the top 15 most influential words were selected for generating explanations for each abstract. We also tried to perform minor adjustments in kernel width or the number of samples, but since this did not significantly alter the outcomes, other parameters were set to their default values.
Figure 3: AOPCglobal for different feature importance metrics. The higher the decrease in predictive performance when top-ranked words are removed (from left to right), the better the feature importance method is in identifying salient features.
Figure 3 shows a summary plot of the AOPCglobal comparison results for LIME, SMER, and a baseline that removed randomly selected words. Table 4 shows the area under the AOPC curve. The random selection, shown by a dashed line, remains almost constant at zero, showing that random word removal has no significant effect on the model’s predictions as expected. An increasing number of removed words results in decreasing changes in the average predicted class probability for both SMER and LIME, which indicates that words associated with the highest scores according to these methods, which were removed first, had the biggest influence. Out of the two aggregation methods for LIME, Global Average LIME Importance had better results; nevertheless, the best overall results were obtained by SMER.
To further validate that this finding extends to other datasets, we compared the performance of SMER and LIME using the same AOPC-based methodology on the IMDB movie sentiment dataset unrelated to the biomedical domain. This evaluation showed the same pattern, with SMER achieving even better results in terms of the area under the AOPC curve compared to LIME. A link to this analysis and results is included in the Availability of code and data section.
Global metric Area under the AOPCglobal curve
SMER SCORE 0.71
Global Average LIME Importance 0.69
Global LIME Importance 0.33
Random baseline 0
Table 4: Area under the AOPCglobal curve for global feature importance metrics. A higher number indicates a better result.
Experiment 4: Qualitative evaluation of top-predicted biological entities
The previous three experiments validated SMER through the quantitative evaluation of aggregate results using standard methodologies: train-test splits (experiment 1) and feature importance metrics (experiments 2 and 3). The goal of this experiment is to validate SMER by evaluating the utility of specific predictions of significant recently introduced virus strains, drugs and vaccines.
To train the model on the most recent data, we used articles published between 2019 and 2020 as the set of training instances for the logistic regression model. The citation counts were used from the year 2021. Word vectors (embeddings) were trained using the embeddings-based model on recent articles published in the years 2019-2022. The rationale for including the year 2022 in the training of the embeddings was to simulate the situation where the text of articles is available but not their actual impact (citation counts).
Table 5 shows the top 15 words sorted by CScore from words that did not appear before the year 2021. These words can be divided into three groups: 1) variants of the SARS-CoV-2 virus, 2) vaccines, and 3) medicines and drugs.
(a) words on positions 1-8 by score
Word topic SCORE NumArt CScore
omicron virus 0.998 1,253 1250
b1351 virus 0.972 981 954
b16172 virus 0.998 714 713
b11529 virus 0.999 349 349
vitt condition 0.844 205 173
b1617 virus 0.998 122 122
casirivimab drug 0.996 92 92
covaxin drug 0.958 95 91
(b) words on position 9-15 by score
Word topic SCORE Count CScore
b16171 virus 0.992 91 90
imdevimab drug 0.997 86 86
pango virus 0.994 84 83
covishield vaccine 0.615 130 80
t478k mutation 0.988 73 72
b1526 virus 1.000 69 69
e484q mutation 0.991 69 68
Table 5: Top 15 words sorted by $CScore=SCORE*NumArt$, where SCORE is the predicted impact using SMER (Eq. 4) and NumArt (Eq. 16) is the count of articles containing the word. Topic is a manually assigned semantic category. Words in bold were selected for a follow-up analysis in Table 6.
From Table 5, representatives of each of the three topics were selected by the highest score value and used as seeds to retrieve additional highly-scored words in each of these three domains. This was done by choosing all words whose word-vector similarity (cosine similarity) was at least 70% to the seed word, sorting the words by SCORE, and taking the first 15 hits and excluding the seed word. The results are shown in Table 6(a) (virus variants — words similar to “b11529” - common SARS-CoV-2 variant), Table 6(b) (drugs — similar to “imdevimab”) and Table 6(c) (vaccines — similar to “covaxin”). In addition to the three tables generated automatically by selecting the highest ranked words in each category, we generated a fourth table (Table 6(d)) containing the highest predicted words by similarity to “Remdesivir”. This choice of the seed word followed a previous analysis [24] that detected topics related to COVID-19 and identified the antiviral drug Remdesivir as the dominant term for the drug topic.
An interesting observation is that Table 6(b) contains several terms related to the Moderna vaccine (’moderna’, ’spikevax’, ’modernas’, ’mRNA-1273’). The two main terms from this list (by frequency) are ”moderna” and ”mRNA-1273”, both with a count of about 800 and very similar importance (SCORE) of 0.9548 and 0.9681, respectively. The high agreement in the score for these close synonyms, which, however, have distinct word vectors and were thus handled as two completely separate words, show that the prediction results are consistent. This also indicates that when the count of articles where the term appears is low, the score estimates are less stable, as can be seen with ”spikevax” (has a count of 41) and ”modernas” (count of only 16), which both have lower scores.
The word “Remdesivir” on its own has with $\textit{SCORE}(w)=0.961$ also a high value of the predicted significance, although not as high as for imdevimab ($\textit{SCORE}(w)=0.997$), which was thus chosen as the representative for the drug topic.
(a) virus strains (“b11529”)
Word SCORE Count Year Sim
b1526 0.9999 69 2021 0.83
vocs 0.9987 810 2004 0.75
g614 0.9984 107 2020 0.71
b16172 0.9981 714 2021 0.90
omicron 0.9979 1,253 2021 0.90
202012/01 0.9977 68 2020 0.82
b1617 0.9977 122 2021 0.86
delta 0.9973 2,873 1996 0.84
b1429 0.9971 59 2021 0.79
b1427 0.9967 33 2021 0.78
ba1 0.9966 91 2020 0.83
d614g 0.9963 867 2020 0.80
501yv2 0.9963 78 2020 0.84
ba2 0.9959 75 2021 0.82
(b) vaccines (“covaxin”)
Word SCORE Count Year Sim
bbv152 0,9810 53 2020 0.77
mrna-1273 0.9681 779 2020 0.72
moderna 0.9548 890 2020 0.73
sinovac 0.9448 109 2020 0.75
bbibp-corv 0.9117 81 2020 0.78
ncov-19 0.8864 544 2020 0.71
mrna1273 0.8818 35 2020 0.71
coronavac 0.8767 302 2020 0.74
gam-covid-vac 0.8715 56 2020 0.77
sputnik 0.8031 143 2020 0.78
chadox1 0.7777 767 2013 0.72
bnt162b2 0.6938 2,210 2020 0.73
spikevax 0.6444 41 2021 0.71
modernas 0.6321 16 2020 0.75
(c) drugs (“imdevimab”)
Word SCORE Count Year Sim
casirivimab 0.9962 92 2021 0.98
etesevimab 0.9929 46 2021 0.91
bamlanivimab 0.9914 179 2020 0.83
sotrovimab 0.9894 48 2021 0.85
regen-cov 0.9882 27 2021 0.84
ly-cov555 0.9669 30 2020 0.81
casirivimab-imdevimab 0.9653 27 2021 0.76
regn-cov2 0.9603 35 2020 0.79
ronapreve 0.9278 8 2021 0.74
regdanvimab 0.8993 16 2021 0.71
monoclonal 0.8918 3,922 1982 0.71
ly-cov016 0.8878 16 2020 0.71
regn10987 0.8676 19 2020 0.71
casirivimab/ imdevimab 0.8332 41 2021 0.84
(d) similar to “remdesivir”
Word SCORE Count Year Sim
arbidol 0.9996 201 2008 0.76
molnupiravir 0.9996 132 2020 0.80
ciclesonide 0.9934 52 2012 0.70
nirmatrelvir 0.9909 27 2022 0.71
lpv/r 0.9896 79 2006 0.74
umifenovir 0.9876 141 2014 0.78
lopinavir-ritonavir 0.9803 141 2020 0.82
paxlovid 0.9756 16 2022 0.75
lopinavir/ritonavir 0.9750 671 2004 0.85
sarilumab 0.9741 121 2019 0.71
ritonavir 0.9714 331 2004 0.79
favipiravir 0.9703 807 2009 0.89
nitazoxanide 0.9522 132 2009 0.79
eidd-2801 0.9298 40 2020 0.75
Table 6: Top words for selected topics. Sorted by the $\textit{SCORE}(w)$, Sim is the cosine similarity of the word vector to the word vector of the topic in the caption of each table, year indicates the earliest publication year in the corpus when the word appears.
Experiment 5: Qualitative evaluation of top-predicted articles
This experiment is based on the same data and predictive model as the previous experiment. However, unlike Experiment 4, which focuses on predictions of entities, this one focuses on the prediction of impactful articles. The predictions are validated through the bibliometric profile of the publishing journals and authors. We analyzed “young” articles published from the year 2021 which have the greatest potential (Score) according to the model.
Figure 4 analyzes a subset of articles with the highest predicted impact (Score), which were published in journals with impact factor. Among the most recent articles (published in 2022), the article that appeared in a journal assigned the highest impact factor value (39.2) is also predicted to have the highest impact according to our method (SMER score). Out of the eight articles with an impact factor, seven appeared in journals assigned an impact factor higher than the median of 4.74 in the source corpus. Note that the impact factor or other journal metadata were not part of the training corpus. Similarly, the authors of most top-ranked articles included an author with an H-index exceeding the median for full professors on journal editorial boards. The high impact factors of journals in which the predicted impactful articles were published, as well as the track-record of their authors, confirm the ability of our model to identify future research impact. The document object identifiers (DOIs) of the specific top-predicted articles are available in the supplementary material.
Figure 4: Top articles by predicted impact (Score) published in 2021. The reference impact factor (blue line) represents the median among the top 10% journals in the category “Health sciences” category. The reference H-index (red line) corresponds to the median H-index of full professors. The attention score and tweets were retrieved from Altmetrics.com.
Discussion
In this section, we present a more detailed examination and discussion of the predicted impactful entities from Experiment 4.
The inspection of the first year of appearance shows that in Table 6(a), there are a few words that predate the beginning of the COVID-19 pandemic in the year 2019. For example, “Delta” is a generic word, which has additional specific meaning in the COVID-19 context where it denotes the name of the virus mutation. Similarly, “ba1” is used to denote a variant of the SARS-CoV-2 virus from the year 2021. For the year 2020, the word “ba1” occurred just once as an antibody unrelated to COVID-19. The oldest word is “monoclonal”, which appeared first in the year 1982 (in the studied corpus). Despite its relative age, it is predicted to have a high impact, possibly related to the fact that monoclonal antibodies are currently (as of 2024) one of the most successful treatment options for COVID-19. We note that all words in this table correspond to specific SARS-CoV-2 variants, with one exception being ’vocs’. This word corresponds to ’variants of concern’ and represents a more general term for variants. Additionally, ’202012/01’ corresponds to another variant of concern ’VOC 202012/01’ [25].
An important property of embeddings is that they can provide predictions even for words that were not contained in the labelled training data for the classification model. This is not possible with models based on the bag of words representation. An example is “nirmatrevil”, which has a high predicted score (0.99) according to our proposed embeddings-based logistic regression model. Since this word only appeared for the first time in an article published in the year 2022, it was not included in the abstracts of any of the articles on which the logistic regression model was trained.
In Table 6(c), we show results for the most important drug-related entities which are similar to the COVID-19 drug imdevimab. To summarize these findings, we note that of the top hits to this drug, seven of these top words are other COVID-19 antibody drugs with the antibody-drug designated suffix -mab. Potentially other methods designed for learning word subtokens could identify further impactful patterns based on word prefix and suffix patterns. However, we observe that the identified important features similar to imdevimab are perfectly categorized as antibody drugs, combinations, or more abstract related terms (’monoclonal’), and include many synonyms. Thus included are: regen-cov (representing the combination of casirivimab (#2) and imdevimab (#1)), ly-cov555 (synonym for bamlanivimab (#4)), regn-cov2 (synonym for casirivimab-imdevimab), ronapreve (synonym for casirivimab-imdevimab), ly-cov016 (also known as etesevimab (#3)), and regn10987 (synonym for imdevimab). In total, including identified synonyms, 8 out of the top important and similar words to imdevimab are synonyms of this drug or its combinations (e.g. the combinations casirivimab-imdevimab, casirivimab-imdevimab, regn-cov2 and ronapreve). The remaining words among the top hits are other COVID-19 antibody drugs, with monoclonal representing the general concept of monoclonal antibodies. With the exception of this general term, all other important words first appeared in the literature in 2020 or later, further validating their relevance for similarity to imdevimab, which is a new drug developed for COVID-19.
In Table 6(d), we also considered similar and important features related to the drug remdevisir, which has been used as a diagnostic word to assess word similarities for COVID-19 literature. In this case, all of the top important features were small-molecule COVID-19 drugs, with one exception being the antibody drug arilumab approved for rheumatoid arthritis and repurposed for COVID-19. We note again that the suffix -ir, indicating an antiviral compound, is found for 8 of the 14 most important words similar to remdesivir, suggesting that subtoken approaches may provide additional benefits. Included in these top hits are some synonyms, e.g. molnupiravir and eidd-2801 (a preclinical name for this drug), lopinavir-ritonavir as well as lopinavir/ritonavir and lpv/r. There are 11/14 instances of first mentions in the literature prior to 2020, all representing drugs that were repurposed for COVID-19. Thus, we note that with our methodology, it is also possible to automatically identify instances of drug repurposing from the literature.
Limitations
The embeddings-based approach to predicting research impact achieves AUC-Extreme10% values of 0.77 for the reference setup. While an AUC in the range between 0.7 and 0.8 is considered as acceptable [26], this does not mean that our method can, with acceptable reliability, assess the research impact of any article. Instead, our research was aimed at evaluating the ability of machine learning models to identify the most promising research.
In our work, we use the widely used Word2vec method to generate embeddings. The newer alternative approaches, such as BERT [27], have the disadvantage that they generate different word embeddings for each context. The use of contextual embeddings might improve predictive performance, but the adoption of stable word embeddings empowers our method to deliver simple explanations. In future work, we plan to extend the method by supporting contextual embeddings, such as those generated by BERT or Large Language Models, such as LLama [28].
In many applications, it is useful for features to correspond to multiword sequences such as bigrams or trigrams. Naturally, by defining $A:=(w,w^{\prime})$ for an entity-pair $w,w^{\prime}$, we can compute $\textit{SCORE}(A)$ using (3). Sorting the entities according to their scores can reveal which connections of words have a high potential to predict the high citability of a paper. Processing $n$-grams can have significance, especially in scientific areas such as medicine or material engineering that often use multi-word expressions to denote concepts.
In our work, we predict whether the paper will belong to the lowly/highly cited category, which may appear to be too coarse-grained compared to directly predicting the citation counts when it comes to identifying the ground-breaking research. However, even in this binary classification framework, the confidence of the prediction to the highly cited class correlates with actual research impact as shown through our experiments. Nevertheless, the methodology could be further extended by directly predicting citation counts. While in our previous work [5], we report on such an approach having lower predictive performance than building a binary classifier as done in the present article, the availability of specific counts of predicted citations could have various applications.
Conclusions
This article presents two interconnected contributions. Our original aim was to devise a supervised framework for predicting which research will become impactful. However, when we started working on this, we did not find a classification approach that would be both embedding-based and directly explainable. As the first contribution, we describe a new feature importance algorithm SMER that can explain a supervised logistic regression model trained on an embeddings-based representation by applying the model to documents corresponding to embedding word vectors for single words. We first theoretically proved that this explanation approach has 100% fidelity with respect to the underlying classification model. Next, we experimentally demonstrated the superiority of embeddings-based models over the bag-of-words representation in terms of predictive performance and the quality of explanations if word importance is calculated with SMER. In another experiment, we showed that SMER produces a better global feature importance ranking than would follow from averaged LIME scores.
These experiments have shown the merits of the method quantitatively, but does it also generate useful predictions? Our second contribution is a new methodology for identifying specific newly published but potentially highly impactful entities as well as complete research articles. We provide a detailed analysis of the significance of the top-predicted drugs, vaccines and viruses. The predictions are also validated through bibliometric analysis, which shows that the method correctly identified research written in some of the best journals by leading scientists. Note that while most scientometric research relies on bibliometric factors such as journal quality, number of authors, etc., our training methodology was based solely on the content of the articles as reflected in the article abstracts, and the bibliometric information is used purely for validation purposes. Therefore our method is less susceptible to introducing institutional, gender, and other biases into the machine learning models.
Declarations
Availability of code and data
We used the implementation of logistic regression from scikit-learn package [29]. As the text data, we used the CORD-19 corpus [11]. The citations dataset is available from the authors at https://github.com/beranoval/PRECOG/blob/main/citationcounts_oci_revised_year.zip.
Code is available from the authors at https://github.com/beranoval/PRECOG/tree/main. The code for Experiment 3 is at https://github.com/adrianakubecova/LIME_DP/blob/main/Lime.ipynb. The supplementary analysis comparing SMER with LIME on the IMDB dataset is at https://github.com/adrianakubecova/LIME_DP/blob/main/Lime_imdb.ipynb.
Competing interests
The authors have no potential conflicts of interest to disclose.
Funding
The work of M. Černý was supported by Czech Science Foundation under project 22-19353S. The work of L. Beranová Dvořáčková was supported by grant IGA 16/2022 “PRECOG: Predicting REsearch COncepts of siGnificance”. T. Kliegr was supported by the long-term support for research activities of FIS VSE. The authors would like to Karel Douda for acquiring the citation counts.
Author contributions
L.D. data processing and analysis, V.S. journal quality data and analysis, A.K. comparison between LIME and SMER and AOPC analysis, M.P.J. analysis targets and interpretation of results, T.K. and L.D. writing, M.Č. formalization, A.D. L.D. and T.K. methodology, T.K. conceptualization, approval of the final manuscript, all authors.